OpenStudy (anonymous):

Evaluate 3x-y-2z if x = 2, y = 1, and z = -2. Now determine two common mistakes that can be made when evaluating expressions and how you can prevent them

OpenStudy (anonymous):

9. Distributing the signs and pemdas.

OpenStudy (anonymous):

Can you help me evaluate it?

OpenStudy (trr54):

3x-y-2z =3(2)-1-2(-2) =6-1+4 =9 Did that make more sense?

OpenStudy (anonymous):

kind of, but could you go through the steps for me, like explain what you did?